OpenStudy (anonymous):

What are the vertex and x-intercepts of the graph of the function below? y =(x + 4)(x - 2)

ganeshie8 (ganeshie8):

to find the x-intercepts, put y=0 and solve x

ganeshie8 (ganeshie8):

y =(x + 4)(x - 2) 0 = (x + 4)(x - 2)

ganeshie8 (ganeshie8):

can u solve x ?

OpenStudy (anonymous):

Vertex: (-1, -9); Intercepts: x = -4, 2

ganeshie8 (ganeshie8):

correct ! you solve it already ?

OpenStudy (anonymous):

yea